# Pseudo-real quantum fields

### v1

@@ 3.1 Issues with the standard approach
The ‘standard’ approach described above has several issues. First, the field $\phi$ cannot be a real scalar field. This can be seen directly from the equations of motion
$\Box_{x}\phi(x)+m^{2}\phi(x)+\frac{ig}{2}\phi(x)^{2}=0\;,$ (3.9a)
$\Box_{x}\phi^{{\dagger}}(x)+m^{2}\phi^{{\dagger}}(x)-\frac{ig}{2}\phi^{{\dagger}2}(x)=0\;,$ (3.9b)
@@
which are not compatible under Hermitian conjugation if we take $\phi=\phi^{{\dagger}}$. Indeed, if we solve these equations perturbatively in $g$, we see that $\phi$ differs from a real field already at order $g$:
$\phi(x)=\phi_{0}(x)+\frac{ig}{2}\int{\rm d}^{4}y\>G_{0}(x,y)\phi_{0}^{2}(y)+\mathcal{O}(g^{2})\>.$ (3.10)
@@
Here, $\phi_{0}$ is the free real Klein-Gordon scalar field and $G_{0}(x,y)$ is the free Klein-Gordon Green’s function, i.e.,
$\quantity(\Box_{x}+m^{2})\phi_{0}(x)$ $=0\;,$ (3.11)
$(\Box_{x}+m^{2})G_{0}(x,y)$ $=-\delta^{(4)}(x-y)\;.$ (3.12)
@@
We can also look at Hamilton’s equations of motion for the field function and the field operator
$\quantity{\phi(x),H}_{\rm PB}=\partial_{t}\phi(x)\text{\>\>\>and\>\>\>}[\hat{\phi}(x),\hat{H}]=i\partial_{t}\hat{\phi}(x)\>,$ (3.13)
@@
where “PB” indicated the Poisson bracket. By taking their Hermitian conjugates, we see that $\phi^{\dagger}$ does not evolve with the same Hamiltonian as $\phi$
$\quantity{\phi^{\dagger}(x),H^{\dagger}}_{\rm PB}=\partial_{t}\phi^{\dagger}(x)\text{\>\>\>and\>\>\>}[\hat{\phi}^{\dagger}(x),\hat{H}^{\dagger}]=i\partial_{t}\hat{\phi}^{\dagger}(x)\>.$ (3.14)
Hence, for any canonical formalism with a non-Hermitian Hamiltonian, the fields must be complex. This observation also shows that the Hermitian-conjugate field $\phi^{\dagger}$ does not transform under the same representation of time translations as $\phi$, which we discuss in the following subsection in the context of the proper Poincaré group. Moreover, the introduction of a complex scalar field doubles the degrees of freedom of the original theory.
@@
The second issue arises from assuming the ‘standard’ parity and time-reversal transformations in Eqs. (3.3a) and (3.3b). Using pseudo-Hermiticity of the Hamiltonian in Eqs. (3.8a) and (3.8b), we find that Hamilton’s equations transform inconsistently:
$P:\>\quantity{\phi(x),H}_{\rm PB}=\partial_{t}\phi(x)$$\longmapsto\quantity{\phi(x),H^{\dagger}}_{\rm PB}=\partial_{t}\phi(x)\>,$ (3.15a)
$T:\>\quantity{\phi(x),H}_{\rm PB}=\partial_{t}\phi(x)$$\longmapsto\quantity{\phi(x),H^{\dagger}}_{\rm PB}=\partial_{t}\phi(x)\>,$ (3.15b)
$\hat{\mathcal{P}}[\hat{\phi}(x),\hat{H}]\hat{\mathcal{P}}^{-1}=i\hat{\mathcal{P}}\partial_{t}\hat{\phi}(x)\hat{\mathcal{P}}^{-1}$ $\longmapsto[\hat{\phi}(x),\hat{H}^{\dagger}]=i\partial_{t}\hat{\phi}(x)\>,$ (3.15c)
$\hat{\mathcal{T}}[\hat{\phi}(x),\hat{H}]\hat{\mathcal{T}}^{-1}=-i\hat{\mathcal{T}}\partial_{t}\hat{\phi}(x)\hat{\mathcal{T}}^{-1}$ $\longmapsto[\hat{\phi}(x),\hat{H}^{\dagger}]=i\partial_{t}\hat{\phi}(x)\>.$ (3.15d)
@@
Hence, the parity and time-reversal transformations which make the theory pseudo-Hermitian and $PT$-symmetric are not the ‘standard’ transformations but instead are non-trivial (see Refs. [17, 18, 16]). We will find them explicitly in the next subsections.
3.2 Pseudo-Hermitian formalism
@@
For a quantum field theory with a non-Hermitian Hamiltonian, the field $\phi$ and its Hermitian conjugate $\phi^{\dagger}$ transform under different representations of the proper Poincaré group. Thus, not only do they not evolve with the same Hamiltonian, but also the Lagrangian composed of $\phi$ and $\phi^{\dagger}$ is not Poincaré invariant [16]. Instead in Ref. [16] (see also Refs. [17, 18, 19]), we propose to build the Lagrangian from the field $\phi$ and its ‘dual’ field $\tilde{\phi}^{\dagger}$, which transform in the same representation of the proper Poincaré group.
@@
Given that the Lagrangian is pseudo-Hermitian with respect to some Hermitian operator $\hat{\eta}$, i.e., $\hat{\eta}\hat{\mathscr{L}}(x_{\eta})\hat{\eta}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)$ and $\hat{\eta}^{\dagger}=\hat{\eta}$, the ‘dual’ field operator $\hat{\tilde{\phi}}^{\dagger}$ has a general form [16]
$\hat{\tilde{\phi}}^{{\dagger}}(x)=\hat{\eta}^{-1}\hat{\phi}^{\dagger}(x_{\eta})\hat{\eta}\>\pi\>.$ (3.16)
@@
Here, $x_{\eta}$ is the coordinate transformation under $\hat{\eta}$ (e.g., parity $x_{P}$). For an $n$-component field $\hat{\phi}^{a}$, where $a\in\{1,\cdots,n\}$, $\pi$ is a Hermitian $n\times n$ matrix. Hence, for a $1$-component scalar field, $\pi\in\mathbb{R}$ is some real constant to be determined.
The ‘dual’ field operator evolves with the same Hamiltonian as $\phi$ and thus has the same equation of motion:
$[\hat{\tilde{\phi}}^{{\dagger}}(x),\hat{H}]=i\partial_{t}\hat{\tilde{\phi}}^{{\dagger}}(x)\>,\qquad\ \Box_{x}\hat{\tilde{\phi}}^{{\dagger}}(x)+m^{2}\hat{\tilde{\phi}}^{{\dagger}}(x)+\frac{ig}{2}\hat{\tilde{\phi}}^{{\dagger}^{2}}(x)=0\>.$ (3.17)
We can also write the ‘dual’ field as a field function in terms of the map $\eta$ (3.4):
$\tilde{\phi}(x)\coloneqq\pi\eta[\phi(x)]\>\implies\>\tilde{\phi}^{\dagger}(x)=\eta^{\dagger}[\phi^{\dagger}(x)]\pi\>,$ (3.18)
which relates the transformation $\eta$ to the operator $\hat{\eta}$ via:
$\eta[\hat{\phi}(x)]=\hat{\eta}\hat{\phi}(x_{\eta})\hat{\eta}^{-1}\>.$ (3.19)
@@
We can take Eq. (3.19) as a definition of the functional map $\eta$. This map is linear, and the Hermiticity of the operator $\hat{\eta}$ implies that the transformation $\eta^{\dagger}=\eta^{-1}$ is unitary, which is consistent with providing a degenerate system of constraints that determine a contour in the complex plane (see Section 2).
@@
Thus, we formulate the $i\phi^{3}$ theory in Eq. (3.1) in terms of a complex scalar field $\phi$, which has a ‘dual’ field $\tilde{\phi}^{\dagger}$. However, a complex scalar field has two propagating degrees of freedom rather than the one degree of freedom described by the original real scalar field. In the following subsection, we argue that the field $\phi$, while being complex, maintains only one degree of freedom. We do this by replacing ‘real’ numbers with ‘pseudo-real’ numbers.
3.3 Pseudo-reality
@@
We propose that, for a pseudo-Hermitian Hamiltonian, ‘real’ fields are not equal to their Hermitian conjugate, i.e., $\phi^{\dagger}\neq\phi$. Instead, they are equal to their ‘dual’ field (3.16), making them ‘pseudo-real’:
$\tilde{\phi}^{\dagger}(x)=\phi(x)\>.$ (3.20)
@@
This constraint is consistent with the Hamiltonian being $\hat{\eta}$-pseudo-Hermitian, i.e., $\hat{H}^{\dagger}=\hat{\eta}\hat{H}\hat{\eta}^{-1}$, and the field $\phi$, while being complex, has the the same number of degrees of freedom as the original real scalar field.
@@
The pseudo-reality condition gives a non-trivial relationship between $\phi$ and its Hermitian conjugate $\phi^{\dagger}$:
$\hat{\phi}^{\dagger}(x)=\pi\hat{\eta}\hat{\phi}(x_{\eta})\hat{\eta}^{-1}=\pi\eta[\hat{\phi}(x)]\>,$ (3.21)
where we have used Eq. (3.19) relating the map $\eta$ to the operator $\hat{\eta}$.
@@
Equation (3.21) presents a consistent resolution of the so-called Hermiticity Puzzle. Previous resolutions of this problem [14] have required a redefinition of Hermitian conjugation, where real scalar fields, such as $\phi$ here, are odd under Hermitian conjugation, i.e., anti-Hermitian. In the present context, we see that this redefinition of Hermitian conjugation amounts to associating it with the composition of Hermitian conjugation and $\eta$-conjugation.
The definition (3.21) of the pseudo-reality of the field $\phi$ is an operator version of the definition (2.4) of the pseudo-reality of a complex number $z$.
3.4 Parity and time-reversal
@@
Let us assume that the field $\phi$ in the non-Hermitian Lagrangian (3.1) is complex and pseudo-real. This Lagrangian still has an anti-linear symmetry and is pseudo-Hermitian, but not with respect to the ‘standard’ parity and time-reversal transformations in Eqs. (3.3a) and (3.3b). Instead, we have a linear ‘parity-like’ and an anti-linear ‘time-reversal-like’ transformation, which map the field $\phi$ to its ‘dual’ conjugate $\tilde{\phi}$. For the ‘parity-like’ transformation $\eta_{P}$, we have:
$\eta_{P}:\>$ $\>\>\phi(x)\longmapsto\phi^{\eta_{P}}(x_{P})=\eta_{P}[\phi(x)]=-\tilde{\phi}(x)\>,\>\>\>\>\>\qquad\qquad\hat{\eta}_{P}\hat{\phi}(x_{P})\hat{\eta}_{P}^{-1}=-\hat{\tilde{\phi}}(x)\>,$ (3.22a)
$\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta_{P}}(x_{P})=\mathscr{L}^{\dagger}(x)\>,\qquad\qquad\qquad\qquad\>\>\>\>\hat{\eta}_{P}\hat{\mathscr{L}}(x_{P})\hat{\eta}_{P}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>,$ (3.22b)
@@
where the constant $\pi_{P}=-1$ in Eq. (3.16) reflects the pseudo-scalar nature of the field. For the ‘time-reversal-like’ transformation $\eta_{T}$, we have:
$\eta_{T}:\>$ $\>\>\phi(x)\longmapsto\phi^{\eta_{T}}(x_{T})=\eta_{T}[\phi(x)]=\tilde{\phi}(x)\>,\qquad\qquad\>\>\>\>\>\>\>\>\hat{\eta}_{T}\hat{\phi}(x_{T})\hat{\eta}_{T}^{-1}=\hat{\tilde{\phi}}(x)\>,$ (3.23a)
$\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta_{T}}(x_{T})=\mathscr{L}^{\dagger}(x)\>,\qquad\qquad\qquad\qquad\>\>\>\>\hat{\eta}_{T}\hat{\mathscr{L}}(x_{T})\hat{\eta}_{T}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>,$ (3.23b)
@@
where the constant is $\pi_{T}=+1$, as the anti-linearity of the time-reversal operator changes the sign of the complex interaction term in the Lagrangian. Hence, the Lagrangian is $\eta_{P}^{-1}\eta_{T}$-symmetric:
$\eta_{P}^{-1}\eta_{T}:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta_{P}^{-1}\eta_{T}}(x_{PT})=\mathscr{L}(x)\>,\>\>\>(\hat{\eta}_{P}^{-1}\hat{\eta}_{T})\hat{\mathscr{L}}(x_{PT})(\hat{\eta}_{P}^{-1}\hat{\eta}_{T})^{-1}=\hat{\mathscr{L}}(x)\>.$ (3.24)
@@
However, in subsection $3.6$, we show that the operators $\hat{\eta}_{P}$ and $\hat{\eta}_{T}$ are related to the ‘standard’ parity $\hat{\mathcal{P}}$ and time-reversal $\hat{\mathcal{T}}$ transformations, such that $\hat{\eta}_{P}^{-1}\hat{\eta}_{T}=\hat{\mathcal{P}}\hat{\mathcal{T}}$, meaning the Lagrangian is still $PT$-symmetric with respect to the ‘standard’ $PT$ transformation.
@@
The pseudo-reality condition in Eq. (3.20) relates the field $\phi$ to its Hermitian conjugate $\phi^{\dagger}$ via the linear operator $\hat{\eta}_{P}$:
$\hat{\phi}^{\dagger}(x)=-\hat{\eta}_{P}\hat{\phi}(x_{P})\hat{\eta}_{P}^{-1}=-\eta_{P}[\hat{\phi}(x)]\>,$ (3.25)
and via the anti-linear operator $\hat{\eta}_{T}$:
$\hat{\phi}^{\dagger}(x)=\hat{\eta}_{T}\hat{\phi}(x_{T})\hat{\eta}_{T}^{-1}=\eta_{T}[\hat{\phi}(x)]\>.$ (3.26)
@@
As we will show in the following subsections, both of these conditions are equivalent and give a single constraint on the field $\phi$.
3.5 Finding $\eta$
In this subsection, we find the explicit form of the transformations $\eta_{P}$ and $\eta_{T}$. We begin with the ‘parity-like’ transformation $\eta_{P}$:
$\eta_{P}:\>\>\phi(x)\longmapsto\eta_{P}[\phi(x)]=-\tilde{\phi}(x)\>,$ (3.27)
where we have used the ‘pseudo-reality’ condition in Eq. (3.20). We see that the field $\eta_{P}[\phi(x)]$ solves the same equation of motion as the field $-\tilde{\phi}(x)$:
$\Box_{x}\eta_{P}[\phi(x)]+m^{2}\eta_{P}[\phi(x)]+\frac{ig}{2}\quantity(\eta_{P}[\phi(x)])^{2}=0\>.$ (3.28)

### v2

@@ III.1 Issues with the standard approach
Despite its apparent consistency, the ‘standard’ approach described above has several issues. First, the field $\phi$ cannot be a real scalar field. This can be seen directly from the equations of motion
$\Box_{x}\phi(x)+m^{2}\phi(x)+\frac{ig}{2}\phi^{2}(x)=0\;,$ (25a)
$\Box_{x}\phi^{{\dagger}}(x)+m^{2}\phi^{{\dagger}}(x)-\frac{ig}{2}\phi^{{\dagger}2}(x)=0\;,$ (25b)
@@
which are not compatible under Hermitian conjugation if we take $\phi=\phi^{{\dagger}}$. Indeed, if we solve these equations perturbatively in $g$, we see that $\phi$ is complex at order $g$:
$\phi(x)=\phi_{0}(x)+\frac{ig}{2}\int{\rm d}^{4}y\>G_{0}(x,y)\phi_{0}^{2}(y)+\mathcal{O}(g^{2})\>,$ (26)
@@
where $\phi_{0}$ is the free real Klein-Gordon scalar field and $G_{0}(x,y)$ is the Green’s function of the free Klein-Gordon operator, i.e.,
$\quantity(\Box_{x}+m^{2})\phi_{0}(x)$ $=0\;,$ (27)
$(\Box_{x}+m^{2})G_{0}(x,y)$ $=-\delta^{(4)}(x-y)\;.$ (28)
@@
We can also look at Hamilton’s equations of motion for the $c$-number field and the field operator:
$\quantity{\phi(x),H}_{\rm PB}=\partial_{t}\phi(x)$  $\text{\>\>\>and\>\>\>}[\hat{\phi}(x),\hat{H}]=i\partial_{t}\hat{\phi}(x)\>,$ (29)
@@
where “PB” is the Poisson bracket. By taking the Hermitian conjugates of Eq. (29), we see that $\phi^{\dagger}$ does not evolve with the same Hamiltonian as $\phi$, since $H^{\dagger}\neq H$, i.e.,
$\quantity{\phi^{\dagger}(x),H^{\dagger}}_{\rm PB}=\partial_{t}\phi^{\dagger}(x)$  $\text{\>\>\>and\>\>\>}[\hat{\phi}^{\dagger}(x),\hat{H}^{\dagger}]=i\partial_{t}\hat{\phi}^{\dagger}(x)\>.$ (30)
Hence, for any canonical formalism with a non-Hermitian Hamiltonian, the fields must be complex. Moreover, for a pseudo-Hermitian theory, the Hamiltonian $\hat{H}$ and its Hermitian conjugate $\hat{H}^{{\dagger}}$ do not commute, since the left- and right-eigenvectors of a non-Hermitian Hamiltonian are distinct. It then follows that the field $\hat{\phi}$ and its Hermitian conjugate $\hat{\phi}^{{\dagger}}$ do not commute at equal times (except at the initial time surface) by virtue of their respective evolutions with $\hat{H}$ and $\hat{H}^{{\dagger}}$.
@@
The second issue arises from assuming the ‘standard’ parity and time-reversal transformations of the field $\phi$ in Eqs. (19a) and (19b). For example, transforming the Hamilton’s equation (29) under parity and using that the Hamiltonian is parity-pseudo-Hermitian (24a), we have
$\hat{\mathcal{P}}\quantity[\hat{\phi}(x),\hat{H}]\hat{\mathcal{P}}^{-1}=\hat{\mathcal{P}}\>i\partial_{t}\hat{\phi}(x)\>\hat{\mathcal{P}}^{-1}\longmapsto\quantity[\hat{\phi}(x),\hat{H}^{\dagger}]=i\partial_{t}\hat{\phi}(x)\>,$ (31)
@@
which leaves us with a contradiction, because, as argued above, $\hat{\phi}$ and $\hat{\phi}^{\dagger}$ do not commute at equal times. The same contradiction is also found under the action of the ‘standard’ time-reversal transformation. Hence, the parity and time-reversal transformations which make the theory pseudo-Hermitian and $PT$ symmetric are not the ‘standard’ transformations but instead are non-trivial (see Refs. Alexandre:2020gah ; Alexandre:2022uns ; Sablevice:2023odu ). We find them explicitly in the next Section III.4.
III.2 Pseudo-Hermitian formalism
@@
For a quantum field theory with a non-Hermitian Hamiltonian, the field $\phi$ and its Hermitian conjugate $\phi^{\dagger}$ transform under different representations of the proper Poincaré group Sablevice:2023odu . Thus, not only do they not evolve with the same Hamiltonian, but also the Lagrangian composed of $\phi$ and $\phi^{\dagger}$ is not Poincaré invariant. Instead, in Ref. Sablevice:2023odu (see also Refs. Alexandre:2020gah ; Alexandre:2022uns ; Chernodub:2021waz ), we proposed to build the Lagrangian from the field $\phi$ and its ‘dual’ field $\tilde{\phi}^{\dagger}$. We define the ‘dual’ field $\tilde{\phi}^{\dagger}$ as the field transforming under the same representation of the proper Poincaré group as $\phi$.
@@
Given a Lagrangian $\hat{\mathscr{L}}(x)$ that is pseudo-Hermitian with respect to some Hermitian operator $\hat{\eta}=\hat{\eta}^{{\dagger}}$, i.e., $\hat{\eta}\hat{\mathscr{L}}(x_{\eta})\hat{\eta}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)$, the ‘dual’ field operator $\hat{\tilde{\phi}}^{\dagger}$, in general, has a form Sablevice:2023odu
$\hat{\tilde{\phi}}^{{\dagger}}(x)=\hat{\eta}^{-1}\hat{\phi}^{\dagger}(x_{\eta})\hat{\eta}\>\pi\>.$ (32)
@@
Here, $x_{\eta}$ is the coordinate transformed under $\hat{\eta}$ (e.g., parity $x_{P}$). For an $n$-component field $\hat{\phi}^{a}$, where $a\in\{1,\cdots,n\}$, $\pi$ is a Hermitian $n\times n$ matrix. Hence, for a $1$-component scalar field, $\pi\in\mathbb{R}$ is some real constant to be determined. We can check that the ‘dual’ field operator in Eq. (32) evolves with the same Hamiltonian as $\hat{\phi}$:
$\quantity[\hat{\tilde{\phi}}^{{\dagger}}(x),\hat{H}]=i\partial_{t}\hat{\tilde{\phi}}^{{\dagger}}(x)\>$ (33)
and, for the Lagrangian (17), it has the same equation of motion as $\hat{\phi}$; namely,
[…]
The corresponding ‘dual’ $\tilde{\phi}^{\dagger}$ of the $c$-number field $\phi$ can be defined via the map $\eta$ (20):
$\tilde{\phi}(x)\coloneqq\pi\eta[\phi(x)]\>\implies\>\tilde{\phi}^{\dagger}(x)=\eta^{\dagger}[\phi^{\dagger}(x)]\pi\>,$ (35)
so that the map $\eta$ is related to the operator $\hat{\eta}$ via
$\eta[\hat{\phi}(x)]=\hat{\eta}\hat{\phi}(x_{\eta})\hat{\eta}^{-1}\>.$ (36)
@@
We can take Eq. (36) as a definition of the functional map $\eta$. This map is linear, and the Hermiticity of the operator $\hat{\eta}$ implies that the transformation satisfies $\eta^{\dagger}=\eta^{-1}$, which is consistent with providing a degenerate system of constraints that determine a contour in the complex plane (see Section II).
@@
Thus, in the remainder of this paper, we study the $i\phi^{3}$ theory in Eq. (17) as a theory of a complex scalar field $\phi$ and its ‘dual’ field $\tilde{\phi}^{\dagger}$. However, a complex scalar field has two propagating degrees of freedom rather than one degree of freedom as described by the original real scalar field. In the following subsection, we argue that the field $\phi$, while being complex, maintains only one degree of freedom. We do this by replacing ‘real’ numbers with ‘pseudo-real’ numbers.
III.3 Pseudo-reality
@@
To this end, we propose that, for a pseudo-Hermitian Hamiltonian, a ‘real’ field is not equal to its Hermitian conjugate, i.e., $\phi^{\dagger}\neq\phi$. Instead, it is equal to its ‘dual’ field (32), making it ‘pseudo-real’; specifically,
$\tilde{\phi}^{\dagger}(x)=\phi(x)\;,\qquad\hat{\tilde{\phi}}^{\dagger}(x)=\hat{\phi}(x)\>.$ (37)
@@
This constraint is consistent with the Hamiltonian being $\hat{\eta}-$pseudo-Hermitian, i.e., $\hat{H}^{\dagger}=\hat{\eta}\hat{H}\hat{\eta}^{-1}$, and the field operator $\hat{\phi}$, while being complex, describes the same number of degrees of freedom as the original real scalar field.
@@
The pseudo-reality condition gives a non-trivial relationship between $\hat{\phi}$ and its Hermitian conjugate $\hat{\phi}^{\dagger}$:
$\hat{\phi}^{\dagger}(x)=\pi\hat{\eta}\hat{\phi}(x_{\eta})\hat{\eta}^{-1}=\pi\eta[\hat{\phi}(x)]\>,$ (38)
where we have used Eq. (36) relating the map $\eta$ to the operator $\hat{\eta}$.
@@
Equation (38) presents a consistent resolution of the so-called Hermiticity Puzzle. Previous resolutions of this problem Mannheim:2018dur have required a redefinition of Hermitian conjugation, where real scalar fields, such as $\phi$ here, are odd under Hermitian conjugation, i.e., anti-Hermitian. In the present context, we see that this redefinition of Hermitian conjugation amounts to associating it with the composition of Hermitian conjugation and $\eta$-conjugation.
III.4 Parity and time-reversal
@@
Let us now assume that the field $\phi$ in the non-Hermitian Lagrangian (17) is complex and pseudo-real. This Lagrangian still has an anti-linear symmetry and is pseudo-Hermitian, but not with respect to the ‘standard’ parity and time-reversal transformations in Eqs. (19a) and (19b). Instead, we have a linear ‘parity-like’ and an anti-linear ‘time-reversal-like’ transformation, which map the field $\phi$ to its ‘dual’ conjugate $\tilde{\phi}$. For the ‘parity-like’ transformation $\eta_{P}$, we have
$\eta_{P}:\>$ $\>\>\phi(x)\longmapsto\phi^{\eta_{P}}(x_{P})=\eta_{P}[\phi(x)]=-\tilde{\phi}(x)\>,\>\>\>\>\>\qquad\qquad\hat{\eta}_{P}\hat{\phi}(x_{P})\hat{\eta}_{P}^{-1}=-\hat{\tilde{\phi}}(x)\>,$ (39a)
$\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta_{P}}(x_{P})=\mathscr{L}^{\dagger}(x)\>,\qquad\qquad\qquad\qquad\>\>\>\>\hat{\eta}_{P}\hat{\mathscr{L}}(x_{P})\hat{\eta}_{P}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>,$ (39b)
@@
where the constant $\pi_{P}=-1$ in Eq. (32) reflects the pseudo-scalar nature of the field. For the ‘time-reversal-like’ transformation $\eta_{T}$, we have
$\eta_{T}:\>$ $\>\>\phi(x)\longmapsto\phi^{\eta_{T}}(x_{T})=\eta_{T}[\phi(x)]=\tilde{\phi}(x)\>,\qquad\qquad\>\>\>\>\>\>\>\>\hat{\eta}_{T}\hat{\phi}(x_{T})\hat{\eta}_{T}^{-1}=\hat{\tilde{\phi}}(x)\>,$ (40a)
$\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta_{T}}(x_{T})=\mathscr{L}^{\dagger}(x)\>,\qquad\qquad\qquad\qquad\>\>\>\>\hat{\eta}_{T}\hat{\mathscr{L}}(x_{T})\hat{\eta}_{T}^{-1}=\hat{\mathscr{L}}^{\dagger}(x)\>,$ (40b)
@@
with $\pi_{T}=+1$, as the anti-linearity of the time-reversal operator changes the sign of the complex interaction term in the Lagrangian. Hence, the Lagrangian is $\eta_{P}^{-1}\eta_{T}$-symmetric:
$\eta_{P}^{-1}\eta_{T}:\>\mathscr{L}(x)\longmapsto\mathscr{L}^{\eta_{P}^{-1}\eta_{T}}(x_{PT})=\mathscr{L}(x)\>,\>\>\>(\hat{\eta}_{P}^{-1}\hat{\eta}_{T})\hat{\mathscr{L}}(x_{PT})(\hat{\eta}_{P}^{-1}\hat{\eta}_{T})^{-1}=\hat{\mathscr{L}}(x)\>.$ (41)
@@
In subsection $3.6$, we will show that the operators $\hat{\eta}_{P}$ and $\hat{\eta}_{T}$ are related to the ‘standard’ parity $\hat{\mathcal{P}}$ and time-reversal $\hat{\mathcal{T}}$ transformations, such that $\hat{\eta}_{P}^{-1}\hat{\eta}_{T}=\hat{\mathcal{P}}\hat{\mathcal{T}}$, meaning the Lagrangian is, in fact, still $PT$ symmetric with respect to the ‘standard’ $PT$ transformation.
@@
The pseudo-reality condition in Eq. (37) relates the field $\hat{\phi}$ to its Hermitian conjugate $\hat{\phi}^{\dagger}$ via the linear operator $\hat{\eta}_{P}$ through
$\hat{\phi}^{\dagger}(x)=-\hat{\eta}_{P}\hat{\phi}(x_{P})\hat{\eta}_{P}^{-1}=-\eta_{P}[\hat{\phi}(x)]\>,$ (42)
and via the anti-linear operator $\hat{\eta}_{T}$ through
$\hat{\phi}^{\dagger}(x)=\hat{\eta}_{T}\hat{\phi}(x_{T})\hat{\eta}_{T}^{-1}=\eta_{T}[\hat{\phi}(x)]\>.$ (43)
@@
As we will show in the following subsections, both of these conditions are equivalent and give a single constraint on the field $\hat{\phi}$.
III.5 Finding $\eta$
To make explicit the constraints provided by the pseudo-reality conditions in Eqs. (42) and (43), we will now determine the explicit forms of the transformations $\eta_{P}$ and $\eta_{T}$. As we will see, in the case where the non-Hermiticity is controlled by a small parameter, this can be achieved by obtaining a differential equation for the transformed fields that resembles the equation of motion for the field.
We begin with the ‘parity-like’ transformation $\eta_{P}$, given in Eq. (39a),
$\eta_{P}:\>\>\phi(x)\longmapsto\eta_{P}[\phi(x)]=-\tilde{\phi}(x)\>.$ (44)
The field $\eta_{P}[\phi(x)]$ solves the same equation of motion (33) as the field $\tilde{\phi}(x)$, i.e.,
$\Box_{x}\eta_{P}[\phi(x)]+m^{2}\eta_{P}[\phi(x)]+\frac{ig}{2}\quantity(\eta_{P}[\phi(x)])^{2}=0\>.$ (45)